xx WaveToy1DF77

WaveToy1DF77

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 amplitude 10 Scope: private01  REAL

11 Description: Amplitude of initial Gaussian wave

11 Range 10              01  Default: 1.0

11 0:*  10

11 bound  10 Scope: private01  KEYWORD

11 Description: Which boundary condition to apply

11 Range 10              01  Default: zero
11 zero  10 Set boundary values to zero
11 radiation 10 Radiation/wave boundary conditions
11 none  10 Don’t apply any condition

11

11 center  10 Scope: private01  REAL

11 Description: x-position of center of Gaussian wave

11 Range 10              01  Default: 0.5
11 *:*  10 Anywhere

11

11 sigma  10 Scope: private01  REAL

11 Description: Width of Gaussian wave

11 Range 10              01  Default: 0.02
11 *:*  10 Anything

11

7 Interfaces

General Implements: wavetoy1d

Inherits: boundary

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11  evolve1d  10                    01  compact 11 0
11                    10  phi  01 description 11 The evolved scalar field
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 3
11                    10                    01 variable type 11 REAL
11  initial  10                    01  compact 11 0
11                    10  timederiv  01 description 11 For initial data
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  dummy  10                    01  compact 11 0
11                    10  phixtemp  01 description 11 Just for creating initial data
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL
11  coords1d  10                    01  compact 11 0
11                    10  x  01 description 11 The 1D x-coordinate
11                    10                    01 dimensions 11 1
11                    10                    01 distribution 11 DEFAULT
11                    10                    01 group type 11 GF
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn CactusExamples/WaveToy1DF77. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 evolve1d[3] coords1d 11 x

11 x  11 x

11

Scheduled Functions CCTKxSTARTUP x wavetoy1df77xstartup x register wavetoy1df77 banner x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxWRAGH x wavetoy1df77xregistercoords x register wavetoy1df77 coordinates x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxBASEGRID x wavetoy1df77xcoord x set up 1d coordinates x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

11

CCTKxINITIAL x wavetoy1df77xinitial x initial data for 1d wave equation x

11 x 11 Language: 11 fortran
11 x 11 Storage: 11 initial
11 x 11 x  11 dummy
11 x 11 Type:  11 function

11

CCTKxEVOL x wavetoy1df77xevolve x evolution of 1d wave equation x

11 x 11 Language: 11 fortran
11 x 11 Sync:  11 evolve1d
11 x 11 Type:  11 function

11

CCTKxEVOL x wavetoy1df77xboundary x apply boundary conditions x

11 x 11 After:  11 wavetoy1df77xevolve
11 x 11 Language: 11 fortran
11 x 11 Type:  11 function

Aliased Functions x

11 Alias Name: xxxxxxx 11 Function Name:
11 WaveToy1DF77xCoord 11 SpatialCoordinates